### v1

# Identifying Network Structure of Linear Dynamical Systems: Observability and Edge Misclassification

Abstract
This work studies the limitations of uniquely identifying a linear network’s topology from partial measurements of its nodes. We show that the set of networks that are consistent with the measurements are related through the nullspace of the observability matrix for the true network. In doing so, we illustrate how potentially many networks are fully consistent with the measurements despite having topologies that are structurally inconsistent with each other, an often neglected consideration in the design of topology inference methods. We then provide an aggregate characterization of the space of possible networks by analytically solving for the most structurally dissimilar network. We find that when observing over 6% of nodes in random network models (e.g., Erdős-Rényi and Watts-Strogatz) the rate of edge misclassification drops to $\sim$ 1%. Extending this discussion, we construct a family of networks that keep measurements $\epsilon$-“close” to each other, and connect the identifiability of these networks to the spectral properties of an augmented observability Gramian.
I Introduction and Motivation
Delineating the causal relationships or connections in networks such as in groups of neurons or social interactions is of significant interest particularly when these connections are inferred from time series measurements of the nodes in the networks [1, 2, 3, 4, 5]. Despite efforts to develop different inference methods to accomplish this task, methods typically yield varying predictions [1]. Specifically, given some time series measurements of the nodal states, each method will infer a network topology that is consistent with the measurements, but often are structurally inconsistent with each other. Moreover, these methods typically assume that every node in the network can be measured and perturbed, an assumption that is often violated in practice for large scale networks where only a subset of the nodes are accessible. The motivation for this work spurs from interests arising in the field of neuroscience to understand the brain’s network structure referred to as the “connectome”, i.e., the connections between neurons or groups of neurons, and how this structure influences its function. Identifying the neural circuits can prove challenging as many variations to these circuits can display the same behavior [6, 7]. Existing systems theory literature studying parameter identifiability typically employs an input-output view [8, 9]. Furthermore, these notions of identifiability have been particularized to linear network models in which the transfer function of these models are studied [10], which has led to frameworks to understand how partially observed nodes impact identifiability of these systems [11, 12]. However, stimulating and measuring all neurons is infeasible and emphasizes the necessity for understanding limitations of partial measurements in a passive measurement setting, a gap this work aims to fill [^1] For results on network topology identification in nonlinear setting, see companion paper [13]..
Figure 1: Given time series measurements of a selection of the nodes in a network, inference methods often predict different network topologies. This paper focuses on describing the set of possible networks consistent with these measurements.
This prompts a few pertinent questions: given a set of time series measurements, can we determine what other networks could have generated these measurements? Moreover, what is the most dissimilar network that could have generated these measurements? The focus of this work is to explore these questions in detail which in turn allows us to better understand the space of possible networks that could be inferred, we summarize this goal in Fig. 1.
This work establishes necessary and sufficient conditions under which two systems will generate identical measurements when unperturbed by external inputs (Section II). In addition, we provide an aggregate characterization of the set of all networks that generate identical measurements, namely the most structurally dissimilar network (i.e., a proxy for how far networks can be from one another) and provide conditions under which this network can be analytically computed (Section III). We extend our analysis to the case where measurements are close but not necessarily identical establishing links to the spectral properties of an augmented observability Gramian (Section IV).
Notation.  We denote matrices ($A$) and vectors ($x$) with uppercase and lowercase letters respectively. $A_{ij}$ denotes the element in the $i$’th row and $j$’th column of $A$. We denote the $\ell_{p}$ norm of a vector with $\|\cdot\|_{p}$. Moreover, we denote the element-wise (Hadamard) and Kronecker product of two matrices by $\odot$ and $\otimes$ respectively. Let $|\cdot|$ and $\text{sgn}(\cdot)$ be the absolute value and sign function respectively, which are applied element-wise. Let $\mathcal{N}(\cdot)$ and $\mathcal{R}(\cdot)$ correspond to the nullspace and range of a matrix. Lastly, let $\text{diag}(\cdot)$ be a diagonal matrix, and $\text{blkdiag}(\cdot)$ be a block diagonal matrix formed from the elements of their arguments.
II Conditions for Identical Measurements
Consider a system that describes the linear dynamics of a network given as
$\Sigma:\begin{cases}\dot{x}(t)=Ax(t)&x(0)=x_{0}\\ y(t)=Cx(t).\end{cases}$ (1)
Here $A\in\mathbb{R}^{n\times n}$ denotes the system matrix and encodes the weighted adjacency matrix of the network. $x\in\mathbb{R}^{n}$ encodes the states of the nodes, and $y\in\mathbb{R}^{p}$ is a linear measurement of the states through an observation matrix $C\in\mathbb{R}^{p\times n}$. We study the conditions under which a perturbed system,
$\tilde{\Sigma}:\begin{cases}\dot{\tilde{x}}(t)=(A+\Delta)\tilde{x}(t)&\tilde{x}(0)=x_{0}\\ \tilde{y}(t)=C\tilde{x}(t),\end{cases}$ (2)
namely one that evolves under $A+\Delta$, where $\Delta$ is a perturbation to the network edge weights, yields identical measurements to the original system. We now make this notion precise.
Definition 1.
We say that the measurements of the original system (1) and the perturbed system (2) are indistinguishable if $\forall x_{0}\in\mathbb{R}^{n}$ $y(t)=\tilde{y}(t)~~\forall t\geq 0$.
Lemma 1.
Measurements of the systems (1) and (2) are indistinguishable, if and only if $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
Proof.
Let $y(t)$ and $\tilde{y}(t)$ denote the measurements of (1) and (2) respectively. Assume that the measurements are indistinguishable, namely, $y(t)=\tilde{y}(t)~~\forall t$. Then we know that $y(t)=Ce^{At}x_{0}=Ce^{(A+\Delta)t}x_{0}=\tilde{y}(t)$. Moreover,
$y(t)-\tilde{y}(t)=C(e^{At}-e^{(A+\Delta)t})x_{0}=0.$ (3)
Since $x_{0}$ is an arbitrary initial condition this is equivalent to requiring
$C(e^{At}-e^{(A+\Delta)t})=\sum_{k=0}^{\infty}C(A^{k}-(A+\Delta)^{k})\frac{t^{k}}{k!}=0.$ (4)
Since this equation holds $\forall t$, for necessity $(\!\!\implies\!\!)$, we examine that at time $t=0$ we must have that the derivatives and function values must be the same
$\frac{d^{j}}{dt^{j}}\left(\sum_{k=0}^{\infty}C(A^{k}-(A+\Delta)^{k})\frac{t^{k}}{k!}\right)\Bigr|_{t=0}=0~~\forall j\in\mathbb{Z}_{\geq 0}$ (5)
$\iff CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}.$
[…]
$\begin{split}y(t)=\sum_{k=0}^{\infty}CA^{k}\frac{x_{0}t^{k}}{k!}=\sum_{k=0}^{\infty}C(A+\Delta)^{k}\frac{x_{0}t^{k}}{k!}=\tilde{y}(t),\end{split}$ (6)
and thus $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}\implies y(t)=\tilde{y}(t)~~\forall t$, which concludes the proof. ∎
Immediately we can establish the following corollary.
Corollary 1.
The measurements are indistinguishable if and only if $CA^{k}=C(A+\Delta)^{k}$ for $k=0,\cdots,n$.
Proof.
This corollary is a direct consequence of the Cayley-Hamilton Theorem. Namely, from Lemma 1, $y(t)=\tilde{y}(t)~~\forall t\iff CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$. Notice that if $CA^{k}=C(A+\Delta)^{k}\text{ for }k=0,\cdots,n$ we have
$\begin{aligned} CA&=C(A+\Delta)\implies C\Delta=0\\ CA^{2}&=CA(A+\Delta)\implies CA\Delta=0\\ &~~\vdots\\ CA^{n}&=CA^{n-1}(A+\Delta)\implies CA^{n-1}\Delta=0.\end{aligned}$
Which is succinctly written as $CA^{k}\Delta=0$ for $k=0,\cdots,n-1$. Applying the Cayley-Hamilton theorem to express any higher powers of $A$ as a linear combination of $A^{k}$ with $k=0,\cdots,n-1$ by induction we may conclude $CA^{k}\Delta=0\quad\forall k\geq 0$. ∎
The result of Lemma 1 together with Corollary 1 can be rewritten in terms of the observability matrix $\mathcal{O}$ of the original system (1) $\Sigma$.
Theorem 1.
Let $\mathcal{O}$ be the observability matrix of the original system, $\Sigma$ given in (1). Measurements will be indistinguishable if and only if $\mathcal{O}\Delta=0$.
Proof.
From Corollary 1 we know that, $y(t)=\tilde{y}(t)~~\forall t\iff CA^{k}=C(A+\Delta)^{k}~~\text{ for }k=0,\cdots,n$. Moreover, for $k=1$ we see that this condition is satisfied if $C\Delta=0$, for $k=2$ we require $CA\Delta=0$, repeated evaluation establishes
$\begin{bmatrix}C^{\top}&(CA)^{\top}&\dots&(CA^{n-1})^{\top}\end{bmatrix}^{\top}\Delta=\mathcal{O}\Delta=0.$ (7)
Which concludes the proof. ∎
Remark 1.
Theorem 1 is a column-wise constraint on the perturbations $\Delta$. This implies that modifications of the outgoing edges of a node are independent from other nodes.
II-A Set of Possible Networks
Theorem 1 establishes that networks in the set, $\mathcal{A}=\{A+\Delta:\mathcal{O}\Delta=0\}$ all generate the same measurements. Now we consider a simple example to examine the cardinality of this set, consider the following network topology and measurement scheme
$A=\begin{bmatrix}1&1&1&0\\ 0&1&1&0\\ 1&0&0&0\\ 0&1&1&1\end{bmatrix},~~C=\begin{bmatrix}1&0&0&0\end{bmatrix}$. (8)
We note that the number of structurally dissimilar networks, i.e., networks that do not share the same underlying unweighted adjacency matrix, is $2^{n^{2}}$. For a network with $n=4$ nodes as in (8) there are $65536$ possible structurally dissimilar networks. We see that for the choice of $A$ and $C$ in (8)
$\mathcal{N}(\mathcal{O})=\text{span}\bigl\{\begin{bmatrix}0&-1&1&0\end{bmatrix}^{\top},\begin{bmatrix}0&0&0&1\end{bmatrix}^{\top}\bigr\}$. (9)
Hence, if we restrict our view to the first column of $A$, we can modify this column to be any of the following structurally dissimilar columns
$\begin{bmatrix}1\\ 1\\ 0\\ 0\end{bmatrix},\begin{bmatrix}1\\ 1\\ 0\\ 1\end{bmatrix},\begin{bmatrix}1\\ 2\\ -1\\ 0\end{bmatrix},\begin{bmatrix}1\\ 2\\ -1\\ 1\end{bmatrix},\begin{bmatrix}1\\ 0\\ 1\\ 1\end{bmatrix}$. (10)
Notice, the edge $A_{11}=1$ is present in all topologies in $\mathcal{A}$, in fact, all of the incoming edges into node 1 ($A_{1j}~\forall j$) cannot be altered, this motivates formalizing this notion.
Definition 2.
Let the columns of $\Phi$ form a basis for $\mathcal{N}(\mathcal{O})$ and $e_{i}$ be a standard basis vector with a $1$ in the $i$’th component. We call an the edge encoded in $A_{ij}$ structurally essential if
$\Phi^{\top}e_{i}=0.$ (11)
Notice that all of the edges encoded in $A_{1j}$ are structurally essential and cannot be altered. In fact, the entire row which corresponds to the evolution of node 1 is structurally essential.
Definition 3.
We call an edge encoded in $A_{ij}$ structurally decoupled from all other edges if $\exists v$ such that
$\Phi v=e_{i}.$ (12)
Moreover, we refer to edges that are not structurally decoupled as being structurally coupled to some other edges.
Notice, the edge encoded in $A_{41}$ was structurally decoupled from all other edges and was not structurally essential, hence we could freely remove or add this edge. The challenge in enumerating through all possible networks amounts to whether the basis of $\mathcal{N}(\mathcal{O})$ allows the structurally coupled edges in the network to be removed independently of one another, for instance, some structurally coupled edges may be only removed or added simultaneously. An explicit example is found in the fourth column of $A$, in this case the edges encoded in $A_{24}$ and $A_{34}$ are either both present or missing. Similarly, (10) shows that there is no way to simultaneously remove both $A_{21}$ and $A_{31}$. In this particular example there are $6^{3}\times 4=864$ structurally dissimilar networks in $\mathcal{A}$, emphasizing the unwieldy nature of enumerating this set. The combinatorial nature of iterating through changes to each edge motivates developing aggregate characterizations of the set $\mathcal{A}$. As a step towards doing so, we now consider a “worst-case” network that is useful in determining how dissimilar network topologies may be.
III Dissimilar Networks
Having established Theorem 1, and given some true network $A$, we would like to understand the structure of the most dissimilar network that is consistent with the measurements made. Towards understanding this worst case network we formulate a minimization program over the nullspace of the observability matrix, specifically, we consider the following optimization program
$\min_{\Delta}$ $\ \|\text{vec}(Z\odot(A+\Delta))\|_{1}$ (13)
s.t. $\ \mathcal{O}\Delta=0.$
Where the matrix $Z\in\mathbb{R}^{n\times n}$ is defined such that $Z_{ij}=1$ if $A_{ij}\neq 0$ and $Z_{ij}=0$ otherwise. By defining $Z$ in this way and applying the element-wise product of $Z$ and $A+\Delta$ we are encouraging that the solution to this optimization problem will encourage removal of edges that exist in $A$. Using an $\ell_{1}$ norm allows us to encourage sparsity effectively setting these edges to zero.
Now, let $\text{rank}(\mathcal{O})=r$ hence, $\dim(\mathcal{N}(\mathcal{O}))=n-r$ and thus, we can define $\Phi\in\mathbb{R}^{n\times(n-r)}$ such that the columns of $\Phi$ form an orthonormal basis for $\mathcal{N}(\mathcal{O})$ such that $\Phi^{\top}\Phi=I_{r}$. Parameterizing $\Delta=\Phi V$ where $V\in\mathbb{R}^{(n-r)\times n}$, we can write the program in (13) as an unconstrained program
$\min_{V}$ $\ \|\text{vec}(A+Z\odot(\Phi V))\|_{1}.$ (14)
Identifying that
$\text{vec}(A+Z\odot(\Phi V))=a+\text{diag}(z)(I_{n}\otimes\Phi)v$ (15)
where $a=\text{vec}(A)\in\mathbb{R}^{n^{2}}$, $v=\text{vec}(V)\in\mathbb{R}^{n(n-r)}$ and $z=\text{vec}(Z)\in\mathbb{R}^{n^{2}}$, we can rewrite (13) as the following equivalent linear program (LP) [14]
$\min_{t,v}$ $\mathbbm{1}^{\top}t$ (16)
s.t. $\ -t\leq a+\Gamma v\leq t.$
We denote $\Gamma\coloneq\text{diag}(z)(I_{n}\otimes\Phi)\in\mathbb{R}^{n^{2}\times n(n-r)}$ for brevity.
Remark 2.
The constraints in (16) can be written as, $t\geq|a+\Gamma v|$, and the optimal solution $(t^{*},v^{*})$ satisfies $t^{*}=|a+\Gamma v^{*}|$. If it did not, we know $\exists s\geq 0~\text{s.t.}~t^{*}-s=|a+\Gamma v^{*}|$, but, $\mathbbm{1}^{\top}(t^{*}-s)<\mathbbm{1}^{\top}t^{*}$ and hence $t^{*}$ would not be optimal.
This program, (16), can be readily solved with LP solvers, however, we now provide conditions under which this problem admits a closed form analytic solution which is advantageous for large scale networks. We have found that it is typical for the solution of (13) to be identical to that of a relaxed $\ell_{2}$ program.
III-A Equivalence of the $\ell_{1}$ and $\ell_{2}$ Programs
Relaxing the program given in (13) to minimizing the $\ell_{2}$ norm, we arrive at the following optimization program
$\min_{V}$ $\ \frac{1}{2}\|\text{vec}(A+Z\odot(\Phi V))\|_{2}^{2}.$ (17)
As this problem is convex, to find the optimal solution it suffices to take the gradient of the objective yielding
$\Phi^{\top}(A+Z\odot(\Phi V^{*}))=0.$ (18)
We note that this is a linear equation when in vectorized form, namely $(I_{n}\otimes\Phi^{\top})(a+\Gamma v^{*})=0$. Given the solution to this matrix equation, $V^{*}$, we now provide a condition under which this solution is also a solution to (13).
Theorem 2.
The solution, $V^{*}$, to (17) is a solution to the original program (13) if
$\Phi^{\top}(Z\odot\text{sgn}(A+Z\odot(\Phi V^{*})))=0.$ (19)
Proof.
Let $v^{*}=\text{vec}(V^{*})$ where $V^{*}$ is a solution to (17). Moreover, assume that $\Phi^{\top}(Z\odot\text{sgn}(A+Z\odot(\Phi V^{*})))=0$. Recall that the programs (13) and (16) are equivalent, hence we show that for an appropriate choice $(v^{*},t^{*},\lambda^{*})$, where $\lambda^{*}=\begin{bmatrix}{\lambda_{1}^{*}}^{\top}&{\lambda_{2}^{*}}^{\top}\end{bmatrix}^{\top}$ is the dual variable associated with the constraint in (16), the Karush-Kuhn-Tucker (KKT) conditions for (16) are satisfied. We select $t^{*}=|a+\Gamma v^{*}|$, and $\lambda_{1,i}^{*}=1$ and $\lambda_{2,i}^{*}=0$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=1$, $\lambda_{1,i}^{*}=0$ and $\lambda_{2,i}^{*}=1$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=-1$, and lastly $\lambda_{1,i}^{*}=1/2=\lambda_{2,i}^{*}$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=0$. With these choices $\lambda_{1}^{*}-\lambda_{2}^{*}=\text{sgn}(a+\Gamma v^{*})$. The Lagrangian of (16) is
[…]
The corresponding KKT conditions are
$\begin{bmatrix}0\\ \mathbbm{1}\end{bmatrix}+\begin{bmatrix}\Gamma^{\top}&-\Gamma^{\top}\\ -I_{n^{2}}&-I_{n^{2}}\end{bmatrix}\lambda^{*}=0$ (21a)
$\begin{bmatrix}\Gamma&-I_{n^{2}}\\ -\Gamma&-I_{n^{2}}\end{bmatrix}\begin{bmatrix}v^{*}\\ t^{*}\end{bmatrix}+\begin{bmatrix}a\\ -a\end{bmatrix}\leq 0$ (21b)
$\lambda^{*}\geq 0$ (21c)
$\text{diag}(\lambda^{*})\left(\begin{bmatrix}\Gamma&-I_{n^{2}}\\ -\Gamma&-I_{n^{2}}\end{bmatrix}\begin{bmatrix}v^{*}\\ t^{*}\end{bmatrix}+\begin{bmatrix}a\\ -a\end{bmatrix}\right)=0.$ (21d)
With $t^{*}$ as chosen, the feasibility condition (21b) is satisfied. Moreover, (21c) is satisfied by our choice of $\lambda^{*}$. It remains to check (21a) and (21d). Notice, from the stationarity of the Lagrangian (21a) we have that $\lambda_{1}^{*}+\lambda_{2}^{*}=\mathbbm{1}$ which is also satisfied from our choice of $\lambda^{*}$. Rewriting (21d) as
$\begin{split}\text{diag}(\lambda_{1}^{*})(a+\Gamma v^{*}-|a+\Gamma v^{*}|)=0\\ \text{diag}(\lambda_{2}^{*})(-(a+\Gamma v^{*})-|a+\Gamma v^{*}|)=0,\end{split}$ (22)
adding and subtracting these equations to each other we have
$\begin{split}\text{diag}(\lambda_{1}^{*}-\lambda_{2}^{*})(a+\Gamma v^{*})=|a+\Gamma v^{*}|\\ \text{diag}(\lambda_{1}^{*}-\lambda_{2}^{*})|a+\Gamma v^{*}|=a+\Gamma v^{*}.\end{split}$ (23)
Since $\lambda_{1}^{*}-\lambda_{2}^{*}=\text{sgn}(a+\Gamma v^{*})$ these equations are satisfied, and thus (21d) is satisfied. Lastly, we must consider the first n(nr)n(n-r) rows of the stationarity of the Lagrangian (21a), namely we must satisfy
$\Gamma^{\top}(\lambda_{1}^{*}-\lambda_{2}^{*})=\Gamma^{\top}\text{sgn}(a+\Gamma v^{*})=0.$ (24)
Which in matrix form reads as
$\Phi^{\top}(Z\odot\text{sgn}(A+Z\odot(\Phi V^{*})))=0,$ (25)
which is satisfied by assumption and concludes the proof.
III-B Illustrative Example
We now revisit example (8). In Fig. 2 we display the network given in (8) (Left) and the corresponding solution to (13) (Right). We see that the influence from the measured node 1 to the unobserved node 4 is removed, moreover, since all edges ($A_{4j}$) into node 4 are structurally decoupled they are entirely removed. However, since nodes 2 and 3 influence the dynamics of node 1, we see that structurally Fig. 2 (Right) essentially permutes the role of nodes 2 and 3 when compared to Fig. 2 (Left). We note here that the structurally essential edges $A_{1j}$ remain present, as expected, in both networks. Moreover, the condition in Theorem 2 is satisfied and hence it suffices to solve (18) to find the most dissimilar network. This example illustrate a few key points, namely, (i) nodes that do not influence the measured nodes can be removed from the network, (ii) the relationship between unmeasured but influencing nodes can be changed by a simple permutation of their labels.
Figure 2: Network described by (8) (Left) and the equivalent maximally dissimilar network (solution to (13)) (Right). Blue dashed circles indicate which nodes are measured.
IV Networks with $\epsilon$-Close measurements
In practice, measurements are associated with some uncertainty due to noise, furthermore, executing the same experiment twice will yield similar yet non-identical measurements. We now begin to characterize how measurements that are “close” together influence the set of possible network topologies that are consistent with them. Moving forward, we assume both $A$ and $A+\Delta$ are Hurwitz so that the corresponding observability Gramians are well defined.
To analyze this setting we consider an augmented system akin to [15, 16]. Specifically, we define
$\bar{A}\coloneq\text{blkdiag}(A,A+\Delta),~~\bar{C}\coloneq\begin{bmatrix}C&-C\end{bmatrix}$, (26)
and the system $\bar{\Sigma}$ as
$\bar{\Sigma}:\begin{cases}\dot{\bar{x}}(t)=\bar{A}\bar{x}(t)&\bar{x}(0)=\begin{bmatrix}x_{0}&x_{0}\end{bmatrix}^{\top}\\ e(t)=\bar{C}\bar{x}(t).\end{cases}$ (27)
Notice, $\|e(t)\|_{2}$ is a measure of the cumulative error between measurements of the original and perturbed systems.
Theorem 3.
Let $\epsilon>0$ and $\bar{W}_{o}$ be the observability Gramian for the augmented system $\bar{\Sigma}$ given in (27), if $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$, then $\|e(t)\|_{2}<\epsilon$.
Proof.
Assume $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$. Notice
$\begin{split}\|e(t)\|_{2}^{2}&=\int_{0}^{\infty}e(t)^{\top}e(t)dt\\ &=\int_{0}^{\infty}\bar{x}(0)^{\top}e^{\bar{A}^{\top}t}\bar{C}^{\top}\bar{C}e^{\bar{A}t}\bar{x}(0)dt=\bar{x}(0)^{\top}\bar{W}_{o}\bar{x}(0).\end{split}$ (28)
By assumption, $\bar{x}(0)^{\top}\bar{W}_{o}\bar{x}(0)<\bar{x}(0)^{\top}\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I\bar{x}(0)=\epsilon^{2}$. Hence, $\|e(t)\|_{2}<\epsilon.$
For future use, we establish the following definition.
Definition 4.
We say that $\bar{W}_{o}$ is a feasible observability Gramian if $\exists\Delta$ such that
$\bar{W}_{o}\bar{A}+\bar{A}^{\top}\bar{W}_{o}=-\bar{C}^{\top}\bar{C}.$ (29)
Analogous to Section III we now consider a similar program to (13) by modifying the constraints. The following program generates the maximally dissimilar network
$\min_{\Delta,\bar{W}_{o}}$ $\ \|\text{vec}(Z\odot(A+\Delta))\|_{1}$ (30)
s.t. $\ \bar{W}_{o}\bar{A}+\bar{A}^{\top}\bar{W}_{o}=-\bar{C}^{\top}\bar{C}$
$\ \bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|_{2}^{2}}I.$
Remark 3.
By optimizing over $\bar{W}_{o}$, this program is nonconvex (bilinear constraint), however, this program can be solved with branch and bound algorithms in YALMIP [17].
Theorem 3 elucidates the relationship between the observability Gramian, $\bar{W}_{o}$, of $\bar{\Sigma}$ (27) and how “close” the measurements of the two systems will be. With this in mind, we can now consider how the set of similar networks is influenced by the spectral properties of $\bar{W}_{o}$.
IV-A Interpretations Based on Observability Gramian
We now consider the implications of (29), which is also a constraint in the optimization program (30). This equation characterizes all possible Gramians $\bar{W}_{o}$ that correspond to networks with $\epsilon$-close measurements. However, given a feasible $\bar{W}_{o}$, there can exist multiple values of $\bar{A}$, which translate to multiple candidate networks via the perturbation $\Delta$. We are interested in further analyzing this relationship. We leverage and extend techniques from [18] to determine what perturbations $\Delta$ are consistent with a given $\bar{W}_{o}$. Specifically, we will explicitly show the dependence of $\Delta$ on the spectra of $\bar{W}_{o}$.
Since $\bar{W}_{o}$ is a symmetric matrix we can express $\bar{W}_{o}=VDV^{\top}$. Denoting $\text{dim}(\mathcal{N}(\bar{W}_{o}))=l$ we have $V=\begin{bmatrix}V_{o}&V_{\bar{o}}\end{bmatrix}$ is a orthogonal matrix with $V_{o}\in\mathbb{R}^{2n\times(2n-l)}$ and $V_{\bar{o}}\in\mathbb{R}^{2n\times l}$ encoding a basis for the observable subspace and the unobservable subspace of $\bar{\Sigma}$ respectively. Here, $D=\text{blkdiag}(\Lambda,0)$ is a diagonal matrix where $\Lambda\in\mathbb{R}^{(2n-l)\times(2n-l)}$ is full rank. Rewriting (29) by pre and post multiplying by $V^{\top}$ and $V$ respectively we arrive at
$DV^{\top}\bar{A}V+V^{\top}\bar{A}^{\top}VD=-(\bar{C}V)^{\top}(\bar{C}V).$ (31)
Identifying that
$V^{\top}\bar{A}V=\begin{bmatrix}\check{A}_{o}&0\\ \check{A}_{21}&\check{A}_{\bar{o}}\end{bmatrix},~~\bar{C}V=\begin{bmatrix}\check{C}_{o}&0\end{bmatrix}$ (32)
will correspond to an observable decomposition of $\bar{\Sigma}$, (31) succinctly reduces to
$\Lambda\check{A}_{o}+\check{A}_{o}^{\top}\Lambda=-\check{C}_{o}^{\top}\check{C}_{o}.$ (33)
Following [18] we can by pre and post multiply by $\Lambda^{-1/2}$ to yield
$\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}+\Lambda^{-1/2}\check{A}_{o}^{\top}\Lambda^{1/2}=-\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}.$ (34)
This implies the symmetric part of $\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}$ is $\frac{1}{2}\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}$. In general we have that
$\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}=-\frac{1}{2}\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}+S,$ (35)
for an arbitrary skew-symmetric matrix $S$. Finally yielding
$\check{A}_{o}=-\frac{1}{2}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}+\Lambda^{-1/2}S\Lambda^{1/2}.$ (36)
Now we have that
$\bar{A}=V\left(\begin{bmatrix}-\frac{1}{2}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}+\Lambda^{-1/2}S\Lambda^{1/2}&0\\ \check{A}_{21}&\check{A}_{\bar{o}}\end{bmatrix}\right)V^{\top}.$ (37)
Which can be equivalently expressed as
$\bar{A}=\text{blkdiag}(A,A+\Delta)=\underbrace{-\frac{1}{2}V_{o}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}V_{o}^{\top}}_{\text{fixed among all networks}}+\\ V_{o}\Lambda^{-1/2}S\Lambda^{1/2}V_{o}^{\top}+V_{\bar{o}}\check{A}_{12}V_{o}^{\top}+V_{\bar{o}}\check{A}_{\bar{o}}V_{\bar{o}}^{\top}.$ (38)
Thus the potentially many solutions, $(S,\check{A}_{21},\check{A}_{\bar{o}})$, to (38) will characterize the set of possible perturbations, $\Delta$, and hence, resulting networks, $A+\Delta$, that generate similar measurements. Moreover, if two inference methods result in two different network hypotheses $A$ and $\tilde{A}$ with different topologies, one could compute the augmented Gramian (i.e., solve (29) with $)\bar{A}=\text{blkdiag}(A,\tilde{A}$)) and solve (38) to find other topologies that fall within the error defined by $A$ and $\tilde{A}$.
We also see from (38) that for a given feasible observability Gramian all networks share a component related to the observable subsystem of $\bar{\Sigma}$ given in (27), and differ by the skew-symmetric matrix $S$ and the two free matrices $\check{A}_{21}$ and $\check{A}_{\bar{o}}$. To illustrate some of the quantitative trends we consider two cases. If $\Lambda_{ii}\ll 1$, the fixed term will be dominant, thus
$\bar{A}\approx-\frac{1}{2}V_{o}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}V_{o}^{\top}.$ (39)
This implies that only the observable term will contribute to the set of possible perturbed networks, which is consistent with our previous intuition from Section II that the observable portions of the networks must be consistent. Similarly we see that if we tolerate significant deviations in measurements, $\Lambda_{ii}\gg 1$ we have that
$\bar{A}\approx V_{o}\Lambda^{-1/2}S\Lambda^{1/2}V_{o}^{\top}+V_{\bar{o}}\check{A}_{12}V_{o}^{\top}+V_{\bar{o}}\check{A}_{\bar{o}}V_{\bar{o}}^{\top}.$ (40)
Implying that there can be significant variability in the network structure due to the free variables influence.
V Simulations
V-A Identical Measurements with Random Networks
To explore the impacts of partial measurement on topology identification we study edge misclassification in Erdős-Rényi and Watts-Strogatz random networks. The choice to study Watts-Strogatz networks arises from their small-world characteristics that are ubiquitous in brain networks [19]. Specifically, we consider networks with $n=100$ nodes with edge presence probability $p=1/6$, and edge flip probability $\beta=1/30$. For the Watts-Strogatz networks we connect each node to $K=3$ of its neighbors. We then sample from a standard normal distribution to select the edge weights. We randomly select which nodes to measure and compute the corresponding maximally dissimilar networks by solving (16). In each case, we compute the percentage of edges that have been flipped (removed/added), we deemed an edge present if the weight was greater $10^{-5}$. We average the results over $100$ trials with different random seeds, the results are indicated in Fig. 3.
Notably, we observe that when measuring under $6$ nodes the random networks are essentially completely unobservable and the most dissimilar networks effectively change the entire network topologies without affecting the measurements. However, after observing greater than $6$ nodes there is a significant decline in the percentage of edges flipped for both random network models. This indicates a phase transition from completely unobservable networks to completely observable networks.
Figure 3: Percentage of edges flipped as more nodes are measured for Erdős-Rényi and Watts-Strogatz random networks.
V-B $\epsilon$-Close Measurements
To explore the implications of the program in (30), we considered the following network and observation scheme
$A=\begin{bmatrix}-3&1&0\\ 0&-3&0\\ 1&0&-3\end{bmatrix},~~C=\begin{bmatrix}1&0&0\end{bmatrix}$. (41)
Moreover, to avoid solving the nonconvex problem (30) we simply fixed $\bar{W}_{o}$ to ensure it was feasible, and removed the constraint $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|_{2}^{2}}I$. In particular, we have that $\sqrt{\lambda_{\max}(\bar{W}_{o})}=0.6487$ and the corresponding maximally dissimilar network was
$A+\Delta=\begin{bmatrix}-2&0&0.1\\ 0&0&0\\ 0.833&0&-2\end{bmatrix}$. (42)
In Fig. 4 we display the observed trajectories, $y(t)$, for the original system, and the most dissimilar system (i.e., solution to (30)), $\tilde{y}(t)$. We selected the initial condition so that $\|x_{0}\|_{2}=1$. As displayed the trajectories remain close over the entire simulation and the error norm was $\|e(t)\|_{2}=0.0357<0.6487\sqrt{2}$. In Fig. 5 we display the corresponding network topologies. We see that the general intuition developed in Section II is violated as node 2, despite influencing the measured node 1, has all connections removed, and node 3 is now compensating by influencing node 1. In Fig. 5 a direct comparison of the most dissimilar networks that produce identical (Right) and close (Middle) is displayed. Allowing the measurements to be “close” yields significant topological variations.
Figure 4: Comparison of the observed trajectories $y(t)$ and $\tilde{y}(t)$ and corresponding error $e(t)$ of the $\epsilon$-close network.
Figure 5: (Left) Original network, (Middle) network that generates similar measurements that is the most dissimilar (solution to (30)), (Right) most dissimilar network that produces identical measurements (solution to (13)). Blue dashed circles indicate which nodes are measured.
VI Conclusions
In this work, we have provided a framework to understand the structural discrepancies that may arise during the inference of a network’s topology. Moreover, these results provide a direct link between the possible networks that are consistent with measurements made and the observability properties of the network in question. With this, we acknowledge that designing a measurement scheme to ensure desirable observability properties a priori is often infeasible, yet we show that random networks can be recovered provided sufficiently many nodes are measured. We anticipate that future work can utilize the results presented to provide a basis for which new inference algorithms can be designed that take into consideration knowledge of the space of possible networks that are consistent with the measurements taken.

### v2

@@
 Abstract
-This work studies the limitations of uniquely identifying a linear network’s topology from partial measurements of its nodes.
-We show that the set of networks that are consistent with the measurements are related through the nullspace of the observability matrix for the true network.
-In doing so, we illustrate how potentially many networks are fully consistent with the measurements despite having topologies that are structurally inconsistent with each other, an often neglected consideration in the design of topology inference methods.
-We then provide an aggregate characterization of the space of possible networks by analytically solving for the most structurally dissimilar network.
-We find that when observing over 6% of nodes in random network models (e.g., Erdős-Rényi and Watts-Strogatz) the rate of edge misclassification drops to $\sim$ 1%.
-Extending this discussion, we construct a family of networks that keep measurements $\epsilon$-“close” to each other, and connect the identifiability of these networks to the spectral properties of an augmented observability Gramian.
+This work studies the limitations of uniquely identifying the structure (i.e., topology) of a networked linear system from partial measurements of its nodal dynamics.
+In general, many networks can be consistent with these measurements; this is a consideration often neglected by standard network inference methods.
+We show that the space of these networks are related through the nullspace of the observability matrix for the true network.
+We establish relevant metrics to investigate this space, including an analytic characterization of the most structurally dissimilar network that can be inferred, as well as the possibility of mis-inferring presence or absence of edges.
+In simulations, we find that when observing over 6% of nodes in random network models (e.g., Erdős-Rényi and Watts-Strogatz), approximately 99% of edges are correctly classified.
+Extending this discussion, we construct a family of networks that keep measurements $\epsilon$-close to each other, and connect the identifiability of these networks to the spectral properties of an augmented observability Gramian.
 I Introduction and Motivation
-Delineating the causal relationships or connections in networks such as in groups of neurons or social interactions is of significant interest particularly when these connections are inferred from time series measurements of the nodes in the networks [1, 2, 3, 4, 5].
-Despite efforts to develop different inference methods to accomplish this task, methods typically yield varying predictions [1].
-Specifically, given some time series measurements of the nodal states, each method will infer a network topology that is consistent with the measurements, but often are structurally inconsistent with each other.
-Moreover, these methods typically assume that every node in the network can be measured and perturbed, an assumption that is often violated in practice for large scale networks where only a subset of the nodes are accessible.
-The motivation for this work spurs from interests arising in the field of neuroscience to understand the brain’s network structure referred to as the “connectome”, i.e., the connections between neurons or groups of neurons, and how this structure influences its function.
-Identifying the neural circuits can prove challenging as many variations to these circuits can display the same behavior [6, 7].
-Existing systems theory literature studying parameter identifiability typically employs an input-output view [8, 9].
-Furthermore, these notions of identifiability have been particularized to linear network models in which the transfer function of these models are studied [10], which has led to frameworks to understand how partially observed nodes impact identifiability of these systems [11, 12].
-However, stimulating and measuring all neurons is infeasible and emphasizes the necessity for understanding limitations of partial measurements in a passive measurement setting, a gap this work aims to fill [^1] For results on network topology identification in nonlinear setting, see companion paper [13]..
-Figure 1: Given time series measurements of a selection of the nodes in a network, inference methods often predict different network topologies.
-This paper focuses on describing the set of possible networks consistent with these measurements.
-This prompts a few pertinent questions: given a set of time series measurements, can we determine what other networks could have generated these measurements?
-Moreover, what is the most dissimilar network that could have generated these measurements?
-The focus of this work is to explore these questions in detail which in turn allows us to better understand the space of possible networks that could be inferred, we summarize this goal in Fig.
-1.
-This work establishes necessary and sufficient conditions under which two systems will generate identical measurements when unperturbed by external inputs (Section II).
-In addition, we provide an aggregate characterization of the set of all networks that generate identical measurements, namely the most structurally dissimilar network (i.e., a proxy for how far networks can be from one another) and provide conditions under which this network can be analytically computed (Section III).
-We extend our analysis to the case where measurements are close but not necessarily identical establishing links to the spectral properties of an augmented observability Gramian (Section IV).
+Inferring causal relationships between nodes in networks (e.g., regions in the brain) from time series measurements is broadly relevant to the engineering and scientific community [17, 4, 15, 19, 6].
+Various inference methods have been proposed to accomplish this task, but they typically yield varying predictions [17].
+Specifically, given some time series measurements of the nodal dynamics, different methods will infer different network structures (i.e., topologies); these are consistent with the measurements but often different from each other.
+Moreover, these methods typically assume that every node in the network can be measured and perturbed, an assumption that is often violated in practice for large networks where only a subset of nodes are accessible.
+This question is particularly relevant for the field of neuroscience, where there is a concerted effort to understand the brain’s network structure (dubbed the connectome) based on time series measurements of brain regions and neurons/neural populations.
+A key question in the field is how the network structure of the brain influences its function.
+However, identifying network structure can prove challenging, as many different structures can exhibit similar behaviors [12, 14].
+This challenge can be partially addressed by systems theory literature.
+The study of parameter identifiability typically employs an input-output view [10, 7]; these have been particularized to linear network models [18], which has led to frameworks to understand how partially observed nodes impact identifiability of these systems [11, 3].
+However, existing work typically makes engineering-centric assumptions, e.g., assuming that we can excite all or most nodes with known perturbations.
+This is clearly infeasible in the neuroscience setting; most of the time even single perturbations are infeasible or prohibitively costly.
+The focus of this paper is to study the problem of network structure identification for linear dynamical systems under passive partial measurements.
+[^1] For results on network structure identification in the nonlinear setting, see companion paper [9]..
+Figure 1: Given time series measurements of a subset of nodes in a network, inference methods often predict different networks and structures.
+This paper describes the set of possible networks consistent with these measurements and provides characterizations of this set.
+In general, as implied above, many networks can be consistent with partial measurements.
+We are interested in identifying all such networks from given measurements (see Fig.
+1).
+We start by establishing necessary and sufficient conditions under which two systems will generate identical measurements when unperturbed by external inputs (Section II).
+We also provide a formulation to solve for the most structurally dissimilar network (i.e., how bad can our inference be) and provide conditions under which this network can be analytically computed (Section III).
+We extend our analysis to the case where measurements are close but not necessarily identical, establishing links to the spectral properties of an augmented observability Gramian (Section IV).
 Notation.
-We denote matrices ($A$) and vectors ($x$) with uppercase and lowercase letters respectively.
-$A_{ij}$ denotes the element in the $i$’th row and $j$’th column of $A$.
-We denote the $\ell_{p}$ norm of a vector with $\|\cdot\|_{p}$.
-Moreover, we denote the element-wise (Hadamard) and Kronecker product of two matrices by $\odot$ and $\otimes$ respectively.
-Let $|\cdot|$ and $\text{sgn}(\cdot)$ be the absolute value and sign function respectively, which are applied element-wise.
+$A_{ij}$ denotes the element in the $i$-th row and $j$-th column of $A$.
+The $\ell_{p}$ norm of a vector is denoted by $\|\cdot\|_{p}$.
+The element-wise (Hadamard) and Kronecker product by $\odot$ and $\otimes$.
+Let $|\cdot|$ and $\text{sgn}(\cdot)$ be the absolute value and sign function which are applied element-wise.
 Let $\mathcal{N}(\cdot)$ and $\mathcal{R}(\cdot)$ correspond to the nullspace and range of a matrix.
-Lastly, let $\text{diag}(\cdot)$ be a diagonal matrix, and $\text{blkdiag}(\cdot)$ be a block diagonal matrix formed from the elements of their arguments.
+Let $\text{diag}(\cdot)$ $\big(\text{blkdiag}(\cdot)\big)$ be a diagonal (block diagonal) matrix formed from the elements of its argument.
 II Conditions for Identical Measurements
-Consider a system that describes the linear dynamics of a network given as
-$\Sigma:\begin{cases}\dot{x}(t)=Ax(t)&x(0)=x_{0}\\ y(t)=Cx(t).\end{cases}$ (1)
-Here $A\in\mathbb{R}^{n\times n}$ denotes the system matrix and encodes the weighted adjacency matrix of the network.
-$x\in\mathbb{R}^{n}$ encodes the states of the nodes, and $y\in\mathbb{R}^{p}$ is a linear measurement of the states through an observation matrix $C\in\mathbb{R}^{p\times n}$.
+Consider the following networked linear system:
+$\Sigma:\begin{cases}\dot{x}(t)=Ax(t),&x(0)=x_{0}\\ y(t)=Cx(t).\end{cases}$ (1)
+Where $A\in\mathbb{R}^{n\times n}$ encodes the weighted adjacency matrix of the network, $x\in\mathbb{R}^{n}$ encodes the states of the nodes, and $y\in\mathbb{R}^{p}$ is a linear measurement of the states through an observation matrix $C\in\mathbb{R}^{p\times n}$.
 We study the conditions under which a perturbed system,
-$\tilde{\Sigma}:\begin{cases}\dot{\tilde{x}}(t)=(A+\Delta)\tilde{x}(t)&\tilde{x}(0)=x_{0}\\ \tilde{y}(t)=C\tilde{x}(t),\end{cases}$ (2)
-namely one that evolves under $A+\Delta$, where $\Delta$ is a perturbation to the network edge weights, yields identical measurements to the original system. We now make this notion precise.
+$\tilde{\Sigma}:\begin{cases}\dot{\tilde{x}}(t)=(A+\Delta)\tilde{x}(t),&\tilde{x}(0)=x_{0}\\ \tilde{y}(t)=C\tilde{x}(t),\end{cases}$ (2)
+yields identical measurements to the original system. Here, the perturbed system evolves under $A+\Delta$, where $\Delta$ is a perturbation to the network edges.
 Definition 1.
-We say that the measurements of the original system (1) and the perturbed system (2) are indistinguishable if $\forall x_{0}\in\mathbb{R}^{n}$ $y(t)=\tilde{y}(t)~~\forall t\geq 0$.
+The measurements of the original system (1) and the perturbed system (2) are indistinguishable if $\forall x_{0}$, $y(t)=\tilde{y}(t)~~\forall t\geq 0$.
 Lemma 1.
-Measurements of the systems (1) and (2) are indistinguishable, if and only if $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
+Measurements of systems (1) and (2) are indistinguishable, if and only if $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
 Proof.
-Let $y(t)$ and $\tilde{y}(t)$ denote the measurements of (1) and (2) respectively.
-Assume that the measurements are indistinguishable, namely, $y(t)=\tilde{y}(t)~~\forall t$.
-Then we know that $y(t)=Ce^{At}x_{0}=Ce^{(A+\Delta)t}x_{0}=\tilde{y}(t)$.
-Moreover,
+Let $y(t)$ and $\tilde{y}(t)$ denote the measurements of (1) and (2) respectively. Assume that the measurements are indistinguishable, i.e., $y(t)=\tilde{y}(t)~~\forall t$. Then,
 $y(t)-\tilde{y}(t)=C(e^{At}-e^{(A+\Delta)t})x_{0}=0.$ (3)
-Since $x_{0}$ is an arbitrary initial condition this is equivalent to requiring
+Since $x_{0}$ is an arbitrary initial condition this is equivalent to
 $C(e^{At}-e^{(A+\Delta)t})=\sum_{k=0}^{\infty}C(A^{k}-(A+\Delta)^{k})\frac{t^{k}}{k!}=0.$ (4)
-Since this equation holds $\forall t$, for necessity $(\!\!\implies\!\!)$, we examine that at time $t=0$ we must have that the derivatives and function values must be the same
+Since this equation holds $\forall t$, for necessity $(\Rightarrow)$, at time $t=0$ the derivatives and function values must be equal:
 $\frac{d^{j}}{dt^{j}}\left(\sum_{k=0}^{\infty}C(A^{k}-(A+\Delta)^{k})\frac{t^{k}}{k!}\right)\Bigr|_{t=0}=0~~\forall j\in\mathbb{Z}_{\geq 0}$ (5)
-$\iff CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}.$
-[…]
+$\Leftrightarrow CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}.$
+Thus we have shown $y(t)=\tilde{y}(t)~~\forall t\Rightarrow CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
+Now to establish sufficiency $(\Leftarrow)$, assume that $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
+Then we know that
 $\begin{split}y(t)=\sum_{k=0}^{\infty}CA^{k}\frac{x_{0}t^{k}}{k!}=\sum_{k=0}^{\infty}C(A+\Delta)^{k}\frac{x_{0}t^{k}}{k!}=\tilde{y}(t),\end{split}$ (6)
-and thus $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}\implies y(t)=\tilde{y}(t)~~\forall t$, which concludes the proof. ∎
-Immediately we can establish the following corollary.
+and thus $CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}\Rightarrow y(t)=\tilde{y}(t)~~\forall t$, which concludes the proof. ∎
+The following corollary shows that we only require equivalence up to the $n$-th power.
 Corollary 1.
@@
 This corollary is a direct consequence of the Cayley-Hamilton Theorem.
-Namely, from Lemma 1, $y(t)=\tilde{y}(t)~~\forall t\iff CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
+From Lemma 1, $y(t)=\tilde{y}(t)~~\forall t\Leftrightarrow CA^{k}=C(A+\Delta)^{k}~~\forall k\in\mathbb{Z}_{\geq 0}$.
 Notice that if $CA^{k}=C(A+\Delta)^{k}\text{ for }k=0,\cdots,n$ we have
-$\begin{aligned} CA&=C(A+\Delta)\implies C\Delta=0\\ CA^{2}&=CA(A+\Delta)\implies CA\Delta=0\\ &~~\vdots\\ CA^{n}&=CA^{n-1}(A+\Delta)\implies CA^{n-1}\Delta=0.\end{aligned}$
-Which is succinctly written as $CA^{k}\Delta=0$ for $k=0,\cdots,n-1$.
-Applying the Cayley-Hamilton theorem to express any higher powers of $A$ as a linear combination of $A^{k}$ with $k=0,\cdots,n-1$ by induction we may conclude $CA^{k}\Delta=0\quad\forall k\geq 0$.
+$\begin{aligned} CA&=C(A+\Delta)\Rightarrow C\Delta=0\\ CA^{2}&=CA(A+\Delta)\Rightarrow CA\Delta=0\\ &~~\vdots\\ CA^{n}&=CA^{n-1}(A+\Delta)\Rightarrow CA^{n-1}\Delta=0.\end{aligned}$
+Hence, $CA^{k}\Delta=0$ for $k=0,\cdots,n-1$.
+For all $k\geq n$, we can apply the Cayley-Hamilton theorem to express $CA^{k}\Delta$ as a linear combination of its lower powers, i.e., $CA^{k}\Delta=\sum_{m=0}^{n-1}\alpha_{m}CA^{m}\Delta$ for some $\alpha_{m}\in\mathbb{R}$.
+By the above analysis, all terms in this sum are equal to zero.
+Thus, we may conclude that $CA^{k}\Delta=0\quad\forall k\geq 0$.
 ∎
-The result of Lemma 1 together with Corollary 1 can be rewritten in terms of the observability matrix $\mathcal{O}$ of the original system (1) $\Sigma$.
+The result of Lemma 1 and Corollary 1 allow us to relate indistinguishability of (1) and (2) to observability of (1).
 Theorem 1.
@@
 Proof.
-From Corollary 1 we know that, $y(t)=\tilde{y}(t)~~\forall t\iff CA^{k}=C(A+\Delta)^{k}~~\text{ for }k=0,\cdots,n$.
-Moreover, for $k=1$ we see that this condition is satisfied if $C\Delta=0$, for $k=2$ we require $CA\Delta=0$, repeated evaluation establishes
-$\begin{bmatrix}C^{\top}&(CA)^{\top}&\dots&(CA^{n-1})^{\top}\end{bmatrix}^{\top}\Delta=\mathcal{O}\Delta=0.$ (7)
+From Corollary 1 we know that, $y(t)=\tilde{y}(t)~~\forall t\Leftrightarrow CA^{k}=C(A+\Delta)^{k}~~\text{ for }k=0,\cdots,n$.
+For $k=1$ we see that this condition is satisfied if $C\Delta=0$; for $k=2$ we require $CA\Delta=0$.
+Repeated evaluation establishes
+$\begin{bmatrix}C^{\top}&(CA)^{\top}&\dots&(CA^{n-1})^{\top}\end{bmatrix}^{\top}\Delta=\mathcal{O}\Delta=0,$ (7)
 which concludes the proof. ∎
@@
 Theorem 1 is a column-wise constraint on the perturbations $\Delta$. This implies that modifications of the outgoing edges of a node are independent from other nodes.
-II-A Set of Possible Networks
-Theorem 1 establishes that networks in the set, $\mathcal{A}=\{A+\Delta:\mathcal{O}\Delta=0\}$ all generate the same measurements.
-Now we consider a simple example to examine the cardinality of this set, consider the following network topology and measurement scheme
+Theorem 1 implies the set of networks that are indistinguishable is $\mathcal{A}=\{A+\Delta:\mathcal{O}\Delta=0\}$. Consider the following example network and measurement scheme:
 $A=\begin{bmatrix}1&1&1&0\\ 0&1&1&0\\ 1&0&0&0\\ 0&1&1&1\end{bmatrix},~~C=\begin{bmatrix}1&0&0&0\end{bmatrix}$. (8)
-We note that the number of structurally dissimilar networks, i.e., networks that do not share the same underlying unweighted adjacency matrix, is $2^{n^{2}}$.
-For a network with $n=4$ nodes as in (8) there are $65536$ possible structurally dissimilar networks.
-We see that for the choice of $A$ and $C$ in (8)
+The number of structurally dissimilar networks, i.e., networks that do not share the same underlying unweighted adjacency matrix, is $2^{n^{2}}$.
+For a network with $n=4$ nodes, as in (8), there are $65536$ possible structurally dissimilar networks.
+For the choice of $A$ and $C$ in (8)
 $\mathcal{N}(\mathcal{O})=\text{span}\bigl\{\begin{bmatrix}0&-1&1&0\end{bmatrix}^{\top},\begin{bmatrix}0&0&0&1\end{bmatrix}^{\top}\bigr\}$. (9)
-Hence, if we restrict our view to the first column of $A$, we can modify this column to be any of the following structurally dissimilar columns
+Consider the first column of $A$, which can be modified by adding a linear combination of vectors in $\mathcal{N}(\mathcal{O})$ into:
 $\begin{bmatrix}1\\ 1\\ 0\\ 0\end{bmatrix},\begin{bmatrix}1\\ 1\\ 0\\ 1\end{bmatrix},\begin{bmatrix}1\\ 2\\ -1\\ 0\end{bmatrix},\begin{bmatrix}1\\ 2\\ -1\\ 1\end{bmatrix},\begin{bmatrix}1\\ 0\\ 1\\ 1\end{bmatrix}$.
 (10)
-Notice, the edge $A_{11}=1$ is present in all topologies in $\mathcal{A}$, in fact, all of the incoming edges into node 1 ($A_{1j}~\forall j$) cannot be altered, this motivates formalizing this notion.
+Other values are possible; we have shown the five that represent all possible structures (i.e., sparsity patterns).
+Notice that edge $A_{11}=1$ is present in all structures in $\mathcal{A}$.
+In fact, all incoming edges into node 1 ($A_{1j}~\forall j$) cannot be altered, motivating a formalization of this phenomenon:
 Definition 2.
-Let the columns of $\Phi$ form a basis for $\mathcal{N}(\mathcal{O})$ and $e_{i}$ be a standard basis vector with a $1$ in the $i$’th component.
-We call an the edge encoded in $A_{ij}$ structurally essential if
+Let the columns of $\Phi$ form a basis for $\mathcal{N}(\mathcal{O})$ and $e_{i}$ be a standard basis vector with $1$ in the $i$-th component. An edge encoded in $A_{ij}$ is structurally essential if
 $\Phi^{\top}e_{i}=0.$ (11)
-Notice that all of the edges encoded in $A_{1j}$ are structurally essential and cannot be altered. In fact, the entire row which corresponds to the evolution of node 1 is structurally essential.
+In our example, all edges in $A_{1j}$ are structurally essential and cannot be altered without changing measured dynamics.
 Definition 3.
-We call an edge encoded in $A_{ij}$ structurally decoupled from all other edges if $\exists v$ such that
+An edge encoded in $A_{ij}$ is structurally decoupled from all other edges if $\exists v$ such that
 $\Phi v=e_{i}.$ (12)
-Moreover, we refer to edges that are not structurally decoupled as being structurally coupled to some other edges.
-Notice, the edge encoded in $A_{41}$ was structurally decoupled from all other edges and was not structurally essential, hence we could freely remove or add this edge.
-The challenge in enumerating through all possible networks amounts to whether the basis of $\mathcal{N}(\mathcal{O})$ allows the structurally coupled edges in the network to be removed independently of one another, for instance, some structurally coupled edges may be only removed or added simultaneously.
-An explicit example is found in the fourth column of $A$, in this case the edges encoded in $A_{24}$ and $A_{34}$ are either both present or missing.
-Similarly, (10) shows that there is no way to simultaneously remove both $A_{21}$ and $A_{31}$.
-In this particular example there are $6^{3}\times 4=864$ structurally dissimilar networks in $\mathcal{A}$, emphasizing the unwieldy nature of enumerating this set.
-The combinatorial nature of iterating through changes to each edge motivates developing aggregate characterizations of the set $\mathcal{A}$.
-As a step towards doing so, we now consider a “worst-case” network that is useful in determining how dissimilar network topologies may be.
+We refer to edges that are not structurally decoupled as being structurally coupled to some other edges.
+Notice that edge $A_{41}$ is structurally decoupled from all other edges; thus, we can freely remove or add this edge.
+Here, enumerating all possible network structures amounts to determining whether the basis of $\mathcal{N}(\mathcal{O})$ allows the structurally coupled edges in the network to be removed independently of one another.
+Some structurally coupled edges may only be removed or added simultaneously.
+For example, edges $A_{24}$ and $A_{34}$ must both be present or both missing; similarly, (10) shows that there is no way to simultaneously remove both $A_{21}$ and $A_{31}$.
+In this particular example, there are $6^{3}\times 4=864$ structurally dissimilar networks in $\mathcal{A}$, emphasizing the unwieldy nature of enumerating this set.
+The combinatorial iterations through the various structures motivates developing aggregate characterizations of $\mathcal{A}$.
 III Dissimilar Networks
-Having established Theorem 1, and given some true network $A$, we would like to understand the structure of the most dissimilar network that is consistent with the measurements made.
-Towards understanding this worst case network we formulate a minimization program over the nullspace of the observability matrix, specifically, we consider the following optimization program
+Given some true network $A$, we would like to understand the structure of the most dissimilar network that is consistent with the measurements made.
+Towards understanding this “worst-case” network we formulate a minimization program over the set $\mathcal{A}$.
+Defining $Z\in\mathbb{R}^{n\times n}$ such that $Z_{ij}=1$ if $A_{ij}\neq 0$ and $Z_{ij}=0$ otherwise, consider:
 $\min_{\Delta}$ $\ \|\text{vec}(Z\odot(A+\Delta))\|_{1}$ (13)
 s.t. $\ \mathcal{O}\Delta=0.$
-Where the matrix $Z\in\mathbb{R}^{n\times n}$ is defined such that $Z_{ij}=1$ if $A_{ij}\neq 0$ and $Z_{ij}=0$ otherwise.
-By defining $Z$ in this way and applying the element-wise product of $Z$ and $A+\Delta$ we are encouraging that the solution to this optimization problem will encourage removal of edges that exist in $A$.
-Using an $\ell_{1}$ norm allows us to encourage sparsity effectively setting these edges to zero.
+The element-wise product of $Z$ and $A+\Delta$ alongside the use of an $\ell_{1}$ norm encourages the solution of (13) to remove edges that exist in $A$.
 Now, let $\text{rank}(\mathcal{O})=r$ hence, $\dim(\mathcal{N}(\mathcal{O}))=n-r$ and thus, we can define $\Phi\in\mathbb{R}^{n\times(n-r)}$ such that the columns of $\Phi$ form an orthonormal basis for $\mathcal{N}(\mathcal{O})$ such that $\Phi^{\top}\Phi=I_{r}$.
@@
 $\text{vec}(A+Z\odot(\Phi V))=a+\text{diag}(z)(I_{n}\otimes\Phi)v$ (15)
-where $a=\text{vec}(A)\in\mathbb{R}^{n^{2}}$, $v=\text{vec}(V)\in\mathbb{R}^{n(n-r)}$ and $z=\text{vec}(Z)\in\mathbb{R}^{n^{2}}$, we can rewrite (13) as the following equivalent linear program (LP) [14]
-$\min_{t,v}$ $\mathbbm{1}^{\top}t$ (16)
+where $a=\text{vec}(A)\in\mathbb{R}^{n^{2}}$, $v=\text{vec}(V)\in\mathbb{R}^{n(n-r)}$ and $z=\text{vec}(Z)\in\mathbb{R}^{n^{2}}$, we can rewrite (13) using the epigraph form of the problem to arrive at a linear program (LP) [5]:
+$\min_{t,v}$ $\mathds{1}^{\top}t$ (16)
 s.t. $\ -t\leq a+\Gamma v\leq t.$
-We denote $\Gamma\coloneq\text{diag}(z)(I_{n}\otimes\Phi)\in\mathbb{R}^{n^{2}\times n(n-r)}$ for brevity.
-Remark 2.
-The constraints in (16) can be written as, $t\geq|a+\Gamma v|$, and the optimal solution $(t^{*},v^{*})$ satisfies $t^{*}=|a+\Gamma v^{*}|$.
-If it did not, we know $\exists s\geq 0~\text{s.t.}~t^{*}-s=|a+\Gamma v^{*}|$, but, $\mathbbm{1}^{\top}(t^{*}-s)<\mathbbm{1}^{\top}t^{*}$ and hence $t^{*}$ would not be optimal.
-This program, (16), can be readily solved with LP solvers, however, we now provide conditions under which this problem admits a closed form analytic solution which is advantageous for large scale networks.
-We have found that it is typical for the solution of (13) to be identical to that of a relaxed $\ell_{2}$ program.
+Here $\Gamma=\text{diag}(z)(I_{n}\otimes\Phi)\in\mathbb{R}^{n^{2}\times n(n-r)}$.
+Problem (16) can be readily solved with LP solvers.
+We now provide conditions under which this problem admits a closed form analytic solution; this provides better intuition about the problem and is also advantageous for the analysis of very large networks (which translate to very large LPs).
 III-A Equivalence of the $\ell_{1}$ and $\ell_{2}$ Programs
-Relaxing the program given in (13) to minimizing the $\ell_{2}$ norm, we arrive at the following optimization program
+Consider the program given in (14); we now relax it to instead minimize the $\ell_{2}$ norm:
 $\min_{V}$ $\ \frac{1}{2}\|\text{vec}(A+Z\odot(\Phi V))\|_{2}^{2}.$ (17)
-As this problem is convex, to find the optimal solution it suffices to take the gradient of the objective yielding
+This problem is convex; to find the optimal solution, take the gradient of the objective and set it to zero:
 $\Phi^{\top}(A+Z\odot(\Phi V^{*}))=0.$ (18)
-We note that this is a linear equation when in vectorized form, namely $(I_{n}\otimes\Phi^{\top})(a+\Gamma v^{*})=0$.
-Given the solution to this matrix equation, $V^{*}$, we now provide a condition under which this solution is also a solution to (13).
+Note that we can vectorize this to yield a linear equation, i.e., $(I_{n}\otimes\Phi^{\top})(a+\Gamma v^{*})=0$.
+Given the solution $V^{*}$ to matrix equation (18), we now provide a condition under which it is also a solution to (13).
 Theorem 2.
-The solution, $V^{*}$, to (17) is a solution to the original program (13) if
+Let $V^{*}$ be a minimizer for (17). It is also a minimizer to the original program (13) if
 $\Phi^{\top}(Z\odot\text{sgn}(A+Z\odot(\Phi V^{*})))=0.$ (19)
 Proof.
-Let $v^{*}=\text{vec}(V^{*})$ where $V^{*}$ is a solution to (17).
-Moreover, assume that $\Phi^{\top}(Z\odot\text{sgn}(A+Z\odot(\Phi V^{*})))=0$.
-Recall that the programs (13) and (16) are equivalent, hence we show that for an appropriate choice $(v^{*},t^{*},\lambda^{*})$, where $\lambda^{*}=\begin{bmatrix}{\lambda_{1}^{*}}^{\top}&{\lambda_{2}^{*}}^{\top}\end{bmatrix}^{\top}$ is the dual variable associated with the constraint in (16), the Karush-Kuhn-Tucker (KKT) conditions for (16) are satisfied.
+Let $v^{*}=\text{vec}(V^{*})$ where $V^{*}$ is a minimizer of (17).
+Assume that $V^{*}$ satsifies (19).
+Recall that the programs (13) and (16) are equivalent.
+We now show that for an appropriate choice $(v^{*},t^{*},\lambda^{*})$, where $\lambda^{*}=\begin{bmatrix}{\lambda_{1}^{*}}^{\top}&{\lambda_{2}^{*}}^{\top}\end{bmatrix}^{\top}$ are the dual variables associated with the constraints in (16), the Karush-Kuhn-Tucker (KKT) conditions for (16) are satisfied.
 We select $t^{*}=|a+\Gamma v^{*}|$, and $\lambda_{1,i}^{*}=1$ and $\lambda_{2,i}^{*}=0$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=1$, $\lambda_{1,i}^{*}=0$ and $\lambda_{2,i}^{*}=1$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=-1$, and lastly $\lambda_{1,i}^{*}=1/2=\lambda_{2,i}^{*}$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=0$.
@@
 The corresponding KKT conditions are
-$\begin{bmatrix}0\\ \mathbbm{1}\end{bmatrix}+\begin{bmatrix}\Gamma^{\top}&-\Gamma^{\top}\\ -I_{n^{2}}&-I_{n^{2}}\end{bmatrix}\lambda^{*}=0$ (21a)
+$\begin{bmatrix}0\\ \mathds{1}\end{bmatrix}+\begin{bmatrix}\Gamma^{\top}&-\Gamma^{\top}\\ -I_{n^{2}}&-I_{n^{2}}\end{bmatrix}\lambda^{*}=0$ (21a)
 $\begin{bmatrix}\Gamma&-I_{n^{2}}\\ -\Gamma&-I_{n^{2}}\end{bmatrix}\begin{bmatrix}v^{*}\\ t^{*}\end{bmatrix}+\begin{bmatrix}a\\ -a\end{bmatrix}\leq 0$ (21b)
@@
 It remains to check (21a) and (21d).
-Notice, from the stationarity of the Lagrangian (21a) we have that $\lambda_{1}^{*}+\lambda_{2}^{*}=\mathbbm{1}$ which is also satisfied from our choice of $\lambda^{*}$.
+Notice, from the stationarity of the Lagrangian (21a) we have that $\lambda_{1}^{*}+\lambda_{2}^{*}=\mathds{1}$ which is also satisfied from our choice of $\lambda^{*}$.
 Rewriting (21d) as
@@
 Since $\lambda_{1}^{*}-\lambda_{2}^{*}=\text{sgn}(a+\Gamma v^{*})$ these equations are satisfied, and thus (21d) is satisfied.
-Lastly, we must consider the first n(nr)n(n-r) rows of the stationarity of the Lagrangian (21a), namely we must satisfy
+Lastly, we must consider the first n(nr)n(n-r) rows of the stationarity of the Lagrangian (21a),
 $\Gamma^{\top}(\lambda_{1}^{*}-\lambda_{2}^{*})=\Gamma^{\top}\text{sgn}(a+\Gamma v^{*})=0.$ (24)
@@
 We now revisit example (8).
-In Fig.
-2 we display the network given in (8) (Left) and the corresponding solution to (13) (Right).
-We see that the influence from the measured node 1 to the unobserved node 4 is removed, moreover, since all edges ($A_{4j}$) into node 4 are structurally decoupled they are entirely removed.
-However, since nodes 2 and 3 influence the dynamics of node 1, we see that structurally Fig.
-2 (Right) essentially permutes the role of nodes 2 and 3 when compared to Fig.
-2 (Left).
-We note here that the structurally essential edges $A_{1j}$ remain present, as expected, in both networks.
-Moreover, the condition in Theorem 2 is satisfied and hence it suffices to solve (18) to find the most dissimilar network.
-This example illustrate a few key points, namely, (i) nodes that do not influence the measured nodes can be removed from the network, (ii) the relationship between unmeasured but influencing nodes can be changed by a simple permutation of their labels.
-Figure 2: Network described by (8) (Left) and the equivalent maximally dissimilar network (solution to (13)) (Right). Blue dashed circles indicate which nodes are measured.
+In the left half of Fig.
+2, we display the original network (i.e., (8)); in the right half of the figure, we display the minimizer to (13), which is the most structurally dissimilar network that still exhibits identical measurements as the original network.
+In this example, the condition in Theorem 2 is satisfied, so it suffices to solve (18) to find the most dissimilar network.
+Not counting self-dynamics, the two networks have only two edges in common.
+Since all edges ($A_{4j}$) into node 4 are structurally decoupled, they are entirely removed in the dissimilar network.
+Notice that since nodes 2 and 3 influence the dynamics of node 1, the dissimilar network essentially permutes the role of nodes 2 and 3 compared to the original network.
+Conversely, structurally essential edges $A_{1j}$ remain present in both networks, as expected.
+This example illustrate a few key points.
+First, edges connected to nodes that do not influence measured nodes can be removed from the network.
+Second, the relationship between unmeasured but influential nodes can be changed by a simple permutation of their labels.
+Figure 2: (Left) Network described by (8). (Right) The most structurally dissimilar network exhibiting identical measurements as the original network. Blue dashed circles indicate measured nodes.
 IV Networks with $\epsilon$-Close Measurements
-In practice, measurements are associated with some uncertainty due to noise, furthermore, executing the same experiment twice will yield similar yet non-identical measurements.
-We now begin to characterize how measurements that are “close” together influence the set of possible network topologies that are consistent with them.
-Moving forward, we assume both $A$ and $A+\Delta$ are Hurwitz so that the corresponding observability Gramians are well defined.
-To analyze this setting we consider an augmented system akin to [15, 16]. Specifically, we define
-$\bar{A}\coloneq\text{blkdiag}(A,A+\Delta),~~\bar{C}\coloneq\begin{bmatrix}C&-C\end{bmatrix}$, (26)
+In practice, measurements are corrupted by noise, and executing the same experiment twice will yield similar yet non-identical measurements.
+We now extend previous analysis to account for this: we characterize the set of networks who yield similar (not necessarily identical) measurements to the original system.
+We assume both $A$ and $A+\Delta$ are Hurwitz [^2] In practice this assumption is necessary as unstable systems would yield exponentially growing trajectories; it is less meaningful to compare the “closeness” of such trajectories.
+so that the corresponding observability Gramians are well-defined.
+We consider an augmented system akin to [16, 1]:
+$\bar{A}=\text{blkdiag}(A,A+\Delta),~~\bar{C}=\begin{bmatrix}C&-C\end{bmatrix}$, (26)
 and the system $\bar{\Sigma}$ as
-$\bar{\Sigma}:\begin{cases}\dot{\bar{x}}(t)=\bar{A}\bar{x}(t)&\bar{x}(0)=\begin{bmatrix}x_{0}&x_{0}\end{bmatrix}^{\top}\\ e(t)=\bar{C}\bar{x}(t).\end{cases}$ (27)
-Notice, $\|e(t)\|_{2}$ is a measure of the cumulative error between measurements of the original and perturbed systems.
+$\bar{\Sigma}:\begin{cases}\dot{\bar{x}}(t)=\bar{A}\bar{x}(t),&\bar{x}(0)=\begin{bmatrix}x_{0}^{\top}&x_{0}^{\top}\end{bmatrix}^{\top}\\ e(t)=\bar{C}\bar{x}(t).\end{cases}$ (27)
+Here, $\|e(t)\|_{2}$ is a measure of the cumulative error between measurements of the original and perturbed systems.
 Theorem 3.
-Let $\epsilon>0$ and $\bar{W}_{o}$ be the observability Gramian for the augmented system $\bar{\Sigma}$ given in (27), if $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$, then $\|e(t)\|_{2}<\epsilon$.
+Let $\epsilon>0$ and $\bar{W}_{o}$ be the observability Gramian for the augmented system $\bar{\Sigma}$ given in (27).
+If $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$, then $\|e(t)\|_{2}<\epsilon$.
 Proof.
-Assume $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$. Notice
+Assume $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$. Then,
 $\begin{split}\|e(t)\|_{2}^{2}&=\int_{0}^{\infty}e(t)^{\top}e(t)dt\\ &=\int_{0}^{\infty}\bar{x}(0)^{\top}e^{\bar{A}^{\top}t}\bar{C}^{\top}\bar{C}e^{\bar{A}t}\bar{x}(0)dt=\bar{x}(0)^{\top}\bar{W}_{o}\bar{x}(0).\end{split}$ (28)
 By assumption, $\bar{x}(0)^{\top}\bar{W}_{o}\bar{x}(0)<\bar{x}(0)^{\top}\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I\bar{x}(0)=\epsilon^{2}$. Hence, $\|e(t)\|_{2}<\epsilon.$
-For future use, we establish the following definition.
+Theorem 3 elucidates the relationship between $\bar{W}_{o}$, the observability Gramian of (27), and how close the measurements of the two systems will be.
+Now, we leverage it to consider the most structurally dissimilar network that produces $\epsilon$-close measurements as the original network.
+This is similar to Section III; however, we no longer require identical measurements.
+We first establish the following definition.
 Definition 4.
@@
 $\bar{W}_{o}\bar{A}+\bar{A}^{\top}\bar{W}_{o}=-\bar{C}^{\top}\bar{C}.$ (29)
-Analogous to Section III we now consider a similar program to (13) by modifying the constraints. The following program generates the maximally dissimilar network
+The following program generates the maximally dissimilar network (analogous to (13)) for $\epsilon$-close measurements:
 $\min_{\Delta,\bar{W}_{o}}$ $\ \|\text{vec}(Z\odot(A+\Delta))\|_{1}$ (30)
-s.t. $\ \bar{W}_{o}\bar{A}+\bar{A}^{\top}\bar{W}_{o}=-\bar{C}^{\top}\bar{C}$
-$\ \bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|_{2}^{2}}I.$
+[…]
 Remark 2.
-By optimizing over $\bar{W}_{o}$, this program is nonconvex (bilinear constraint), however, this program can be solved with branch and bound algorithms in YALMIP [17].
-Theorem 3 elucidates the relationship between the observability Gramian, $\bar{W}_{o}$, of $\bar{\Sigma}$ (27) and how “close” the measurements of the two systems will be.
-With this in mind, we can now consider how the set of similar networks is influenced by the spectral properties of $\bar{W}_{o}$.
-IV-A Interpretations Based on Observability Gramian
-We now consider the implications of (29), which is also a constraint in the optimization program (30).
-This equation characterizes all possible Gramians $\bar{W}_{o}$ that correspond to networks with $\epsilon$-close measurements.
-However, given a feasible $\bar{W}_{o}$, there can exist multiple values of $\bar{A}$, which translate to multiple candidate networks via the perturbation $\Delta$.
-We are interested in further analyzing this relationship.
-We leverage and extend techniques from [18] to determine what perturbations $\Delta$ are consistent with a given $\bar{W}_{o}$.
-Specifically, we will explicitly show the dependence of $\Delta$ on the spectra of $\bar{W}_{o}$.
-Since $\bar{W}_{o}$ is a symmetric matrix we can express $\bar{W}_{o}=VDV^{\top}$.
-Denoting $\text{dim}(\mathcal{N}(\bar{W}_{o}))=l$ we have $V=\begin{bmatrix}V_{o}&V_{\bar{o}}\end{bmatrix}$ is a orthogonal matrix with $V_{o}\in\mathbb{R}^{2n\times(2n-l)}$ and $V_{\bar{o}}\in\mathbb{R}^{2n\times l}$ encoding a basis for the observable subspace and the unobservable subspace of $\bar{\Sigma}$ respectively.
-Here, $D=\text{blkdiag}(\Lambda,0)$ is a diagonal matrix where $\Lambda\in\mathbb{R}^{(2n-l)\times(2n-l)}$ is full rank.
-Rewriting (29) by pre and post multiplying by $V^{\top}$ and $V$ respectively we arrive at
+This program is nonconvex due to its bilinear constraint, but can be solved with branch and bound algorithms in YALMIP [13].
+We now consider the implications of (29).
+This equation characterizes all possible Gramians $\bar{W}_{o}$ that correspond to networks with $\epsilon$-close measurements to the original.
+However, given a feasible $\bar{W}_{o}$ there can exist multiple values of $\bar{A}$, which translate to multiple candidate networks via the perturbation $\Delta$.
+We extend techniques from [8] to determine what perturbations $\Delta$ are consistent with a given $\bar{W}_{o}$.
+Since $\bar{W}_{o}$ is a real symmetric matrix, we can diagonalize it using an orthonormal matrix $V$: $\bar{W}_{o}=VDV^{\top}$.
+We can partition $V$ as $V=\begin{bmatrix}V_{o}&V_{\bar{o}}\end{bmatrix}$, where $l:=\text{dim}(\mathcal{N}(\bar{W}_{o}))$, $V_{o}\in\mathbb{R}^{2n\times(2n-l)}$, and $V_{\bar{o}}\in\mathbb{R}^{2n\times l}$.
+The columns of $V_{o}$ and $V_{\bar{o}}$ form a basis for the observable and the unobservable subspace of $\bar{\Sigma}$, respectively.
+Diagonal matrix $D$ can be partitioned as $D=\text{blkdiag}(\Lambda,0)$ where $\Lambda\in\mathbb{R}^{(2n-l)\times(2n-l)}$ is full rank.
+We now pre- and post-multiply (29) by $V^{\top}$ and $V$:
 $DV^{\top}\bar{A}V+V^{\top}\bar{A}^{\top}VD=-(\bar{C}V)^{\top}(\bar{C}V).$ (31)
-Identifying that
+Here,
 $V^{\top}\bar{A}V=\begin{bmatrix}\check{A}_{o}&0\\ \check{A}_{21}&\check{A}_{\bar{o}}\end{bmatrix},~~\bar{C}V=\begin{bmatrix}\check{C}_{o}&0\end{bmatrix}$ (32)
-will correspond to an observable decomposition of $\bar{\Sigma}$, (31) succinctly reduces to
+correspond to an observability decomposition of $\bar{\Sigma}$; so (31) reduces to
 $\Lambda\check{A}_{o}+\check{A}_{o}^{\top}\Lambda=-\check{C}_{o}^{\top}\check{C}_{o}.$ (33)
-Following [18] we can by pre and post multiply by $\Lambda^{-1/2}$ to yield
+Following [8], pre- and post-multiply by $\Lambda^{-1/2}$ to obtain
 $\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}+\Lambda^{-1/2}\check{A}_{o}^{\top}\Lambda^{1/2}=-\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}.$ (34)
-This implies the symmetric part of $\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}$ is $\frac{1}{2}\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}$. In general we have that
+This implies that the symmetric part of $\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}$ is $\frac{1}{2}\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}$. In general, we have that
 $\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}=-\frac{1}{2}\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}+S,$ (35)
-for an arbitrary skew-symmetric matrix $S$. Finally yielding
+for an arbitrary skew-symmetric matrix $S$. Pre- and post-multiply again to yield
 $\check{A}_{o}=-\frac{1}{2}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}+\Lambda^{-1/2}S\Lambda^{1/2}.$ (36)
-Now we have that
-$\bar{A}=V\left(\begin{bmatrix}-\frac{1}{2}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}+\Lambda^{-1/2}S\Lambda^{1/2}&0\\ \check{A}_{21}&\check{A}_{\bar{o}}\end{bmatrix}\right)V^{\top}.$ (37)
+Returning to the full augmented system, we have
+$\bar{A}=V\left(\begin{bmatrix}-\frac{1}{2}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}+\Lambda^{-1/2}S\Lambda^{1/2}&0\\ \check{A}_{21}&\check{A}_{\bar{o}}\end{bmatrix}\right)V^{\top},$ (37)
 which can be equivalently expressed as
 $\bar{A}=\text{blkdiag}(A,A+\Delta)=\underbrace{-\frac{1}{2}V_{o}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}V_{o}^{\top}}_{\text{fixed among all networks}}+\\ V_{o}\Lambda^{-1/2}S\Lambda^{1/2}V_{o}^{\top}+V_{\bar{o}}\check{A}_{12}V_{o}^{\top}+V_{\bar{o}}\check{A}_{\bar{o}}V_{\bar{o}}^{\top}.$ (38)
-Thus the potentially many solutions, $(S,\check{A}_{21},\check{A}_{\bar{o}})$, to (38) will characterize the set of possible perturbations, $\Delta$, and hence, resulting networks, $A+\Delta$, that generate similar measurements.
-Moreover, if two inference methods result in two different network hypotheses $A$ and $\tilde{A}$ with different topologies, one could compute the augmented Gramian (i.e., solve (29) with $)\bar{A}=\text{blkdiag}(A,\tilde{A}$)) and solve (38) to find other topologies that fall within the error defined by $A$ and $\tilde{A}$.
-We also see from (38) that for a given feasible observability Gramian all networks share a component related to the observable subsystem of $\bar{\Sigma}$ given in (27), and differ by the skew-symmetric matrix $S$ and the two free matrices $\check{A}_{21}$ and $\check{A}_{\bar{o}}$.
-To illustrate some of the quantitative trends we consider two cases.
+From this analysis, we see that solutions $(S,\check{A}_{21},\check{A}_{\bar{o}})$ to (38) will characterize perturbations $\Delta$ and resulting networks $A+\Delta$ that generate similar measurements.
+We also see from (38) that for a given feasible observability Gramian, all networks share a component related to the observable subsystem of $\bar{\Sigma}$ given in (27), and differ by the skew-symmetric matrix $S$ and the two free matrices $\check{A}_{21}$ and $\check{A}_{\bar{o}}$.
+To illustrate some of the quantitative trends, we consider two cases.
 If $\Lambda_{ii}\ll 1$, the fixed term will be dominant, thus
@@
 This implies that only the observable term will contribute to the set of possible perturbed networks, which is consistent with our previous intuition from Section II that the observable portions of the networks must be consistent.
-Similarly we see that if we tolerate significant deviations in measurements, $\Lambda_{ii}\gg 1$ we have that
-$\bar{A}\approx V_{o}\Lambda^{-1/2}S\Lambda^{1/2}V_{o}^{\top}+V_{\bar{o}}\check{A}_{12}V_{o}^{\top}+V_{\bar{o}}\check{A}_{\bar{o}}V_{\bar{o}}^{\top}.$ (40)
-Implying that there can be significant variability in the network structure due to the free variables influence.
+Similarly we see that if we tolerate significant deviations in measurements, $\Lambda_{ii}\gg 1$ we have
+$\bar{A}\approx V_{o}\Lambda^{-1/2}S\Lambda^{1/2}V_{o}^{\top}+V_{\bar{o}}\check{A}_{12}V_{o}^{\top}+V_{\bar{o}}\check{A}_{\bar{o}}V_{\bar{o}}^{\top},$ (40)
+which implies that there can be significant variability in the network structure due to the free variables’ influence.
 V Simulations
 V-A Identical Measurements with Random Networks
-To explore the impacts of partial measurement on topology identification we study edge misclassification in Erdős-Rényi and Watts-Strogatz random networks.
-The choice to study Watts-Strogatz networks arises from their small-world characteristics that are ubiquitous in brain networks [19].
-Specifically, we consider networks with $n=100$ nodes with edge presence probability $p=1/6$, and edge flip probability $\beta=1/30$.
-For the Watts-Strogatz networks we connect each node to $K=3$ of its neighbors.
-We then sample from a standard normal distribution to select the edge weights.
-We randomly select which nodes to measure and compute the corresponding maximally dissimilar networks by solving (16).
-In each case, we compute the percentage of edges that have been flipped (removed/added), we deemed an edge present if the weight was greater $10^{-5}$.
-We average the results over $100$ trials with different random seeds, the results are indicated in Fig.
+To explore the impact of partial measurement on network structure identification, we study edge misclassification in Erdős-Rényi and Watts-Strogatz random networks; the latter exhibits properties similar to brain networks [2].
+We consider networks with $n=100$ nodes with edge presence probability $p=1/6$ and edge flip probability $\beta=1/30$.
+For the Watts-Strogatz networks, we connect each node to $K=3$ of its neighbors.
+We sample from a standard normal distribution to select edge weights, and randomly select which nodes to measure.
+We compute the maximally dissimilar networks (that exhibit identical dynamics) by solving (16).
+In each case, we compute the percentage of edges that have been flipped (removed/added) relative to the original network; we deemed an edge present if the weight was greater $10^{-5}$.
+Results were averaged over $100$ trials with different random seeds and are depicted in Fig.
 3.
-Notably, we observe that when measuring under $6$ nodes the random networks are essentially completely unobservable and the most dissimilar networks effectively change the entire network topologies without affecting the measurements.
-However, after observing greater than $6$ nodes there is a significant decline in the percentage of edges flipped for both random network models.
-This indicates a phase transition from completely unobservable networks to completely observable networks.
+When measuring under $6$ nodes, the most dissimilar networks effectively change the entire network structure without affecting the measurements.
+However, after observing greater than $6$ nodes, there is a significant decline in the percentage of edges flipped for both network models.
+This indicates a phase transition from completely “unobservable” networks to completely “observable” networks.
 Figure 3: Percentage of edges flipped as more nodes are measured for Erdős-Rényi and Watts-Strogatz random networks.
 V-B $\epsilon$-Close Measurements
-To explore the implications of the program in (30), we considered the following network and observation scheme
+To explore the implications of the program in (30), we consider the following network and observation scheme
 $A=\begin{bmatrix}-3&1&0\\ 0&-3&0\\ 1&0&-3\end{bmatrix},~~C=\begin{bmatrix}1&0&0\end{bmatrix}$. (41)
-Moreover, to avoid solving the nonconvex problem (30) we simply fixed $\bar{W}_{o}$ to ensure it was feasible, and removed the constraint $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|_{2}^{2}}I$.
-In particular, we have that $\sqrt{\lambda_{\max}(\bar{W}_{o})}=0.6487$ and the corresponding maximally dissimilar network was
+For example’s sake, we seek an exact solution; thus, instead of solving nonconvex problem (30), we work backwards.
+We fix $\bar{W}_{o}$ to ensure feasibility and temporarily remove the other constraint.
+Then, we find that $\sqrt{\lambda_{\max}(\bar{W}_{o})}=0.65$ and the corresponding maximally structurally dissimilar network is
 $A+\Delta=\begin{bmatrix}-2&0&0.1\\ 0&0&0\\ 0.833&0&-2\end{bmatrix}$. (42)
 In Fig.
-4 we display the observed trajectories, $y(t)$, for the original system, and the most dissimilar system (i.e., solution to (30)), $\tilde{y}(t)$.
+4, we show the observed trajectories $y(t)$ and $\tilde{y}(t)$ for the original system and most dissimilar system, respectively.
 We selected the initial condition so that $\|x_{0}\|_{2}=1$.
-As displayed the trajectories remain close over the entire simulation and the error norm was $\|e(t)\|_{2}=0.0357<0.6487\sqrt{2}$.
+We see that the trajectories remain close over the entire simulation; the error norm is $\|e(t)\|_{2}=0.036$.
+This is less than $0.65\sqrt{2}$, which is the upper bound provided by Theorem 3.
 In Fig.
-5 we display the corresponding network topologies.
-We see that the general intuition developed in Section II is violated as node 2, despite influencing the measured node 1, has all connections removed, and node 3 is now compensating by influencing node 1.
-In Fig.
-5 a direct comparison of the most dissimilar networks that produce identical (Right) and close (Middle) is displayed.
-Allowing the measurements to be “close” yields significant topological variations.
+5, we display the corresponding network structures.
+We see that the general intuition developed in Section II is violated as node 2, despite influencing measured node 1, has all connections removed; this is compensated by node 3’s influence on node 1.
+For comparison, we also include the most dissimilar network that produces identical measurements; we see that allowing the measurements to be close instead of identical yields significant structural variations.
 Figure 4: Comparison of the observed trajectories $y(t)$ and $\tilde{y}(t)$ and corresponding error $e(t)$ of the $\epsilon$-close network.
-Figure 5: (Left) Original network, (Middle) network that generates similar measurements that is the most dissimilar (solution to (30)), (Right) most dissimilar network that produces identical measurements (solution to (13)).
+Figure 5: (Left) Original network.
+(Middle) The most structurally dissimilar network that generates similar measurements.
+(Right) The most structurally dissimilar network that generates identical measurements.
 Blue dashed circles indicate which nodes are measured.
 VI Conclusions
-In this work, we have provided a framework to understand the structural discrepancies that may arise during the inference of a network’s topology.
-Moreover, these results provide a direct link between the possible networks that are consistent with measurements made and the observability properties of the network in question.
-With this, we acknowledge that designing a measurement scheme to ensure desirable observability properties a priori is often infeasible, yet we show that random networks can be recovered provided sufficiently many nodes are measured.
-We anticipate that future work can utilize the results presented to provide a basis for which new inference algorithms can be designed that take into consideration knowledge of the space of possible networks that are consistent with the measurements taken.
+In this work, we provided tools to understand the structural discrepancies that may arise during network inference from passive partial measurements.
+These results provide a direct link between the set of networks that are consistent with observed measurements and the observability properties of the network in question.
+Future work will utilize the presented results to design new inference algorithms that produce the set of all possible data-consistent networks as opposed to a single possible network.